# Steering Dialogue Dynamics for Robustness against Multi-turn Jailbreaking Attacks

### v1

@@ Abstract
Large language models (LLMs) are highly vulnerable to jailbreaking attacks, wherein adversarial prompts are designed to elicit harmful responses. While existing defenses effectively mitigate single-turn attacks by detecting and filtering unsafe inputs, they fail against multi-turn jailbreaks that exploit contextual drift over multiple interactions, gradually leading LLMs away from safe behavior. To address this challenge, we propose a safety steering framework grounded in safe control theory, ensuring invariant safety in multi-turn dialogues. Our approach models the dialogue with LLMs using state-space representations and introduces a novel neural barrier function (NBF) to detect and filter harmful queries emerging from evolving contexts proactively. Our method achieves invariant safety at each turn of dialogue by learning a safety predictor that accounts for adversarial queries, preventing potential context drift toward jailbreaks. Extensive experiments under multiple LLMs show that our NBF-based safety steering outperforms safety alignment baselines, offering stronger defenses against multi-turn jailbreaks while maintaining a better trade-off between safety and helpfulness under different multi-turn jailbreak methods.
@@
keywords:
AI safety, safe control, neural barrier function
@@ 1 Introduction
Despite the tremendous potential of large language models (LLMs) across a variety of applications, frontier models remain vulnerable to jailbreaking attacks, wherein adversarial prompts are designed to elicit harmful responses (Wei et al., 2023; Anwar et al., 2024; Sun et al., 2024). These attacks include optimization-based methods (Zou et al., 2023; Geisler et al., 2024; Andriushchenko et al., 2024) and automated techniques in which attackers use LLMs to produce jailbreaks (Chao et al., 2023; Liu et al., 2024c; Mehrotra et al., 2023). To counter these threats, defenses such as fine-tuning-based algorithms (Yuan et al., 2024; Zou et al., 2024), inference-time interventions (Arditi et al., 2024; Bhattacharjee et al., 2024; Robey et al., 2023; Li et al., 2024a), and reasoning-based guardrails (Kang and Li, 2024; Liu et al., 2025; Zaremba et al., 2025) have been proposed. And indeed, these defenses have been successful in reducing the effectiveness of single-turn attacks, wherein an adversary can only attempt to jailbreak a targeted model via a single round of conversation.
@@
Unfortunately, the success of current defenses against single-turn jailbreaks has not extended to the more sophisticated setting of multi-turn jailbreaking, wherein an attacker attempts to elicit harmful content throughout multiple rounds of conversation (Li et al., 2024b; Pavlova et al., 2024). In the multi-turn setting, attackers exploit the gradual shift in context to bypass safeguards (Ren et al., 2024; Russinovich et al., 2024; Jiang et al., 2024), making detection and mitigation significantly more challenging, as shown in Figure 1. Multi-turn jailbreaks use adaptive and dynamic interactions to subtly steer LLMs toward unsafe outputs (Zhou et al., 2024; Liu et al., 2024b), circumventing traditional defenses that operate on individual prompts. Therefore, few defenses have been proposed for the multi-turn setting, and it remains unclear how such attacks can be mitigated.
In this work, we approach the problem of defending against multi-turn jailbreaks through the lens of dynamical systems. We view the context of each dialogue turn as an evolving hidden state and define a successful attack as one that transitions into an unsafe region of the state space. Inspired by forward invariance in safe control theory (Liu and Tomizuka, 2014; Ames et al., 2019; Robey et al., 2020), we introduce the concept of invariant safety for dialogue dynamics, ensuring that at every step of a conversation, the system remains within a user-specified safe set, thereby preventing attackers from gradually leading LLMs to generate harmful responses. As shown in Figure 2, we first learn state-space representations corresponding to the neural dialogue dynamics from multi-turn conversations. Then, given these representations, we train a safety predictor as a neural barrier function (NBF) to predict whether the states corresponding to a multi-turn conversation drift into harmful regions of the state space. During the evaluation phase, the NBF filters out potentially harmful queries based on the predictor’s outputs. Extensive experiments validate the effectiveness of the proposed NBF-based steering compared to baselines and show strong generalizability to different LLMs and multi-turn jailbreaking attacks. Our code is available on https://github.com/HanjiangHu/NBF-LLM. In summary, the contributions are listed below.
@@
  • Comprehensive experiments show that the proposed NBF-based safety steering can outperform safety alignment baselines with a better trade-off of safety and helpfulness on multiple LLMs.
@@ 2 Related Work
Jailbreak attacks on LLMs have advanced from hand-crafted prompts to automated red-teaming approaches. Optimization-based methods, including gradient-based and evolutionary attacks (Zou et al., 2023; Geisler et al., 2024; Liu et al., 2023a; Andriushchenko et al., 2024), generate adversarial inputs, while automated attackers leverage LLMs for iterative refinements (Chao et al., 2023; Liu et al., 2024c; Robey et al., 2024). While existing safety alignment (Yuan et al., 2024; Zou et al., 2024), inference-time steering methods (Arditi et al., 2024; Bhattacharjee et al., 2024) and reasoning-based LLM guardrails (Kang and Li, 2024; Liu et al., 2025) are effective against various single-turn jailbreaks, they struggle to defend against multi-turn jailbreaks (Wang et al., 2024; Tong et al., 2024). Multi-turn jailbreaking scenarios are developed by embedding malicious intent gradually (Jiang et al., 2024), breaking down harmful prompts into benign sub-queries (Yu et al., 2024; Zhou et al., 2024; Liu et al., 2024b), designing attack patterns (Ren et al., 2024), and dynamically adjusting attack queries based on contextual feedback (Li et al., 2024b; Yang et al., 2024; Russinovich et al., 2024). Since all these methods gradually steer conversations toward harmful topics, such dynamically drifting context poses a serious challenge to existing reactive safety defense mechanisms. To the best of our knowledge, we are the first to safeguard LLM dialogues dynamically from jailbreaks.
@@
In the literature of control and robotics, there is extensive research on learning-based controllers for dynamical systems that provide safety guarantees or certificates (Boffi et al., 2021; Herbert et al., 2021; Xiao et al., 2023; Lindemann et al., 2021; Chang et al., 2019; Mazouz et al., 2022). Neural networks have been employed to parameterize control barrier functions (CBFs) to achieve forward invariance (Robey et al., 2020; So et al., 2023; Zinage et al., 2023; Dawson et al., 2022; Dai et al., 2022): Once the system states enter the user-defined safe set, they remain within it indefinitely, thereby guaranteeing safety with neural certificates. Although neural CBFs can be successfully learned and verified for control-affined dynamical systems (Manda et al., 2024; Hu et al., 2024b; Mathiesen et al., 2022; Wang et al., 2023; Rickard et al., 2025), it is still challenging to guarantee the safety for non-analytical dynamical systems in latent space (Hu et al., 2024a; Wei and Liu, 2022; Liu et al., 2023b; Li et al., 2023; Shen et al., 2024; Cheng et al., 2024), as the problem we tackle here. Recently, Bajcsy and Fisac (2024) formulate a human-AI safety filter framework and Miyaoka and Inoue (2024) introduce CBFs for LLM safety at the token level against single-turn jailbreaks. However, no existing work has explored dialogue-level safety for LLMs from the perspective of neural CBFs.
@@ 3 Problem Formulation
In this section, we formalize our approach to defending against multi-turn jailbreaks by enforcing invariant safety in the conversation with LLMs. Our framework models the conversation as an evolving dynamical system, where the hidden state represents the drifting context (e.g. the production process in Figure 1), and each dialogue turn represents a state transition influenced by user queries and LLM responses in the language space. Given the language space ${\mathcal{S}}$, at each turn $k$, LLMs receive a query $U_{k}\in{\mathcal{S}}$ from users and make a response $Z_{k}\in{\mathcal{S}}$ to that query, followed by the next round of user query $U_{k+1}\in{\mathcal{S}}$ and LLMs response $Z_{k+1}\in{\mathcal{S}}$. In accordance with the AI usage policies (OpenAI, 2022), LLM responses $Z_{k},k-1,2,\dots,K$ should follow the AI safety rules and fall into the safe region specified by the user ${\mathcal{S}}_{0}$. However, in the multi-turn setting, the context drift of the query along the dynamic conversation may increase the vulnerability of the LLMs for jailbreaks, even though the response is safe at each turn. To this end, we introduce the concept invariant safety and associated measures to guarantee that once the LLM response falls into a safety region – which needs to be computed, the following LLM responses will stay within it no matter what future queries are given along the query context flow.
@@
For any safe but non-invariant responses $Z^{\prime}_{k}\in{\mathcal{S}}_{0}\setminus{\mathcal{S}}_{I}$ (e.g. the LLM response at turn 3 in Figure 1), there exists a potentially harmful query $U^{\prime}_{k+1}\in{\mathcal{S}}_{context}^{(k)}$ such that the next LLM response $Z^{\prime}_{k+1}$ will inevitably go out of ${\mathcal{S}}_{0}$, resulting in LLM jailbreaks. Therefore, the safety invariance subset ${\mathcal{S}}_{I}$ is introduced to avoid non-invariant responses to achieve invariant safety against multi-turn jailbreaking attacks. In the following section, we first model the neural dialogue dynamics of multi-turn conversation with LLMs using the state-space representations. Then, we introduce the safety predictor based neural barrier function, followed by an invariant safety certificate learning framework.
@@ 4.1 Neural Dialogue Dynamics of Multi-turn Conversation with LLMs
Different from studying large language models (LLMs) dynamics of token generation in the literature (Soatto et al., 2023; Liu et al., 2024a; Kong et al., 2024; Miyaoka and Inoue, 2024), we focus on the multi-turn human-AI interactive sentence-wise dynamics in the sentence embedding space and formulate it based on the state space representations in control theory. Given a $K$-turn conversation dialog with user query sentences $U_{k}\in{\mathcal{S}},k=1,\dots K,$ and response sentences of LLMs $Z_{k}\in{\mathcal{S}},k=1,\dots K,$ we first map them from the sentence language space ${\mathcal{S}}$ to the semantic and meaningful embedding space $\mathbb{R}^{n}$ using the pre-trained sentence embedding model $f_{embedding}:{\mathcal{S}}\rightarrow\mathbb{R}^{n}$ (Reimers and Gurevych, 2019; Fonseca et al., 2025) as follows,
@@
We assume the dialogue with LLM is governed by the discrete-time state-transition equation parameterized by neural networks (NNs) $f_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ under initial state $\textbf{0}_{m}\in\mathbb{R}^{m}$, where the user query embedding $u_{k}$ serves as the control input at each turn $k$. However, the state representation $x_{k}$ is partially observable from the response embedding $z_{k}$ because the multi-turn dialogue dynamics is non-Markov due to the memory mechanism of LLMs. Therefore, we formulate the LLM response generation process through another NN-parameterized observation function $g_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where the LLM response embedding $z_{k}$ is the observed output from the hidden state $x_{k}\in\mathbb{R}^{m}$ and the user query embedding $u_{k}$. Assuming embeddings in Equation 2 and states are consistent across different LLM seeds given fixed user queries and temperature, we have,
@@
To learn the state-space representation from the multi-turn dialogue, we construct the following mean square error (MSE) loss given $N$ trajectories of $K$-turn query and response embeddings $u^{(i)}_{k},z^{(i)}_{k}\in\mathbb{R}^{n},i=1,\dots,N,k=1,\dots,K$ from the pre-trained embedding model in Equation 2,
@@ 4.2 Neural Barrier Function based on Safety Predictor
On top of the language state-space dynamics above, we introduce the safety property in this section. Following the literature using LLM-based judge scores to evaluate the performance (Ren et al., 2024; Qi et al., 2023; Zheng et al., 2023), we assume that for each conversation trajectory, the query and response embeddings $u_{k},z_{k}$ at $k$-th turn are associated with a discrete safety score $y_{k}\in{\mathcal{Y}}$ as a label from an LLM judge (Qi et al., 2023). The scores of non-jailbreaking turns fall into the safe subset ${\mathcal{Y}}_{safe}\subset{\mathcal{Y}}$, which is equivalent to the user-specified safety region, i.e. $y_{k}\in{\mathcal{Y}}_{safe}\Leftrightarrow Z_{k}\in{\mathcal{S}}_{0}$. Using a more granular set of safety labels instead of simple binary ones allows for a more nuanced assessment of safety levels, providing richer information for training and evaluation. We adopt a NN parameterized safety predictor $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ to output the difference of classification probability $p(\hat{y}_{k}\mid x_{k-1},u_{k})$ between the safe predicted label and the most unsafe predicted label. Inspired by Cohen et al. (2019), the safety predictor is formally defined as,
@@
where the predicted label can be found through classification model $\hat{y}_{k}=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}p(y\mid x_{k-1},u_{k})$. It can be trained through the cross-entropy loss with $N$ trajectories of $K$-turn queries and responses and state-space dynamics in Equation 3 ,
@@
Now we consider the state evolution in Equation 3 during the multi-turn conversation by bridging the safe control theory (Liu and Tomizuka, 2014; Ames et al., 2014) and the safety predictor. Since the user query sequences determine the state trajectory during the interactive conversation with LLMs, we first denote the query context embedding set at turn $k$ as ${\mathcal{U}}_{k}$, which evolves along the query context flow ${\mathcal{S}}^{(k)}_{context}$ due to multi-turn jailbreaking attacks. To prevent multi-turn jailbreaking, it suffices to ensure that the safety predictor always has safe predictions (negative outputs) given the previous state $x_{k}$ and all potential query embedding $u$ among query context embedding set ${\mathcal{U}}_{k}$ at each turn $k$. Formally, we define the neural barrier function (NBF) below as a safety index to show if the current state can be jailbroken or not along the conversation.
@@
Given the safety predictor $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ defined in Equation 5, denote the query context embedding set at turn $k$ as ${\mathcal{U}}_{k-1}:=\{u\subset\mathbb{R}^{n}\mid u=f_{embedding}(U),\forall U\in{\mathcal{S}}_{context}^{(k-1)}\}$, and then the neural barrier function $\phi_{k}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ and the induced safe set ${\mathcal{X}}_{k}\subset\mathbb{R}^{m}$ are defined as,
@@
where state $x$ follows Equation 3 and $\eta\geq 0$ is the steering threshold w.r.t the safe set ${\mathcal{X}}_{k}$.
Based on the NBF $\phi_{k}$, the safe set ${\mathcal{X}}_{k}$ is defined as the zero sublevel set of $\phi_{k}$ under the context embedding set ${\mathcal{U}}_{k-1}$ at each turn $k$. The larger the steering threshold $\eta$ is, the more shrinkage the induced safe set ${\mathcal{X}}_{k}$ will have. Therefore, $x_{k-1}\in{\mathcal{X}}_{k}$ indicates that the LLM cannot be jailbroken by any potential query $u$ given the conversation query context set ${\mathcal{U}}_{k-1}$, meaning the original safe response sentence $Z_{k-1}$ is within the safety invariance set, i.e., $Z_{k-1}\in{\mathcal{S}}_{I}$. The conversation in the language and embedding space is illustrated in Figure 3.

### v2

@@ Abstract
Large language models (LLMs) are shown to be vulnerable to jailbreaking attacks where adversarial prompts are designed to elicit harmful responses. While existing defenses effectively mitigate single-turn attacks by detecting and filtering unsafe inputs, they fail against multi-turn jailbreaks that exploit contextual drift over multiple interactions, gradually leading LLMs away from safe behavior. To address this challenge, we propose a safety steering framework grounded in safe control theory, ensuring invariant safety in multi-turn dialogues. Our approach models the dialogue with LLMs using state-space representations and introduces a novel neural barrier function (NBF) to detect and filter harmful queries emerging from evolving contexts proactively. Our method achieves invariant safety at each turn of dialogue by learning a safety predictor that accounts for adversarial queries, preventing potential context drift toward jailbreaks. Extensive experiments under multiple LLMs show that our NBF-based safety steering outperforms safety alignment, prompt-based steering and lightweight LLM guardrails baselines, offering stronger defenses against multi-turn jailbreaks while maintaining a better trade-off among safety, helpfulness and over-refusal. Check out the website here https://sites.google.com/view/llm-nbf/home. Our code is available on https://github.com/HanjiangHu/NBF-LLM.
@@ 1 Introduction
Despite the tremendous potential of large language models (LLMs) across a variety of applications, frontier models remain vulnerable to jailbreaking attacks, wherein adversarial prompts are designed to elicit harmful responses (wei2023jailbroken, ; anwar2024foundational, ; sun2024trustllm, ). These attacks include optimization-based methods (zou2023universal, ; geisler2024attacking, ; andriushchenko2024jailbreaking, ) and automated techniques in which attackers use LLMs to produce jailbreaks (chao2023jailbreaking, ; liu2024autodan, ; mehrotra2023tree, ). To counter these threats, defenses such as fine-tuning-based algorithms (yuan2024refuse, ; zou2024improving, ), inference-time interventions (arditi2024refusal, ; bhattacharjee2024towards, ; robey2023smoothllm, ; li2024inference, ), and reasoning-based guardrails (kang2024r, ; liu2025guardreasoner, ; zaremba2025trading, ) have been proposed. These defenses have been successful in reducing the effectiveness of single-turn attacks, wherein an adversary can only attempt to jailbreak a targeted model via a single round of conversation.
@@
Unfortunately, the success of current defenses against single-turn jailbreaks has not extended to the more sophisticated setting of multi-turn jailbreaking, wherein an attacker attempts to elicit harmful content throughout multiple rounds of conversation (li2024llm, ; pavlova2024automated, ). In the multi-turn setting, attackers exploit the gradual shift in context to bypass safeguards (ren2024derail, ; russinovich2024great, ; jiang2024red, ), making detection and mitigation significantly more challenging, as shown in Figure˜1. Multi-turn jailbreaks use adaptive and dynamic interactions to subtly steer LLMs toward unsafe outputs (zhou2024speak, ; liu2024imposter, ), circumventing traditional single-turn defenses. One concurrent defense method (lu2025x, ) proposes to learn more fine-grained boundary-safe and harmful representations through safety fine-tuning, but the multi-turn dialogue context is not explicitly considered and the fine-tuning method is not agnostic to different LLMs as a versatile guardrail (markov2023holistic, ; zeng2024shieldgemma, ; inan2023llama, ).
In this work, we approach the problem of defending against multi-turn jailbreaks through the lens of dynamical systems. We view the context of each dialogue turn as an evolving hidden state and define a successful attack as one that transitions into an unsafe region of the state space. Inspired by forward invariance in safe control theory (liu2014control, ; ames2019control, ; robey2020learning, ), we introduce the concept of invariant safety for dialogue dynamics, ensuring that at every step of a conversation, the system remains within a user-specified safe set, thereby preventing attackers from gradually leading LLMs to generate harmful responses. As shown in Figure˜2, we first learn state-space representations corresponding to the neural dialogue dynamics from multi-turn conversations. Then, given these representations, we train a safety predictor as a neural barrier function (NBF) to predict whether the states corresponding to a multi-turn conversation drift into harmful regions of the state space. During the evaluation phase, the NBF filters out potentially harmful queries based on the predictor’s outputs. Extensive experiments validate the effectiveness of the proposed NBF-based steering and show strong generalizability to different LLMs and multi-turn jailbreaking attacks. In summary, the contributions are listed below.
@@
  • Comprehensive experiments show that the proposed NBF-based safety steering can outperform defense baselines with a better trade-off of safety, helpfulness and over-refusal on multiple LLMs.
@@ 2 Related Work
Jailbreak attacks on LLMs have advanced from hand-crafted prompts to automated red-teaming approaches. Optimization-based methods, including gradient-based and evolutionary attacks (zou2023universal, ; geisler2024attacking, ; liu2023autodan, ; andriushchenko2024jailbreaking, ), generate adversarial inputs, while automated attackers leverage LLMs for iterative refinements (chao2023jailbreaking, ; liu2024autodan, ; robey2024jailbreaking, ). While existing safety alignment (yuan2024refuse, ; zou2024improving, ; zhang2025safety, ), inference-time steering methods (arditi2024refusal, ; bhattacharjee2024towards, ) and reasoning-based LLM guardrails (kang2024r, ; liu2025guardreasoner, ) are effective against various single-turn jailbreaks, they struggle to defend against multi-turn jailbreaks (wang2024mrj, ; tong2024securing, ). Multi-turn jailbreaking scenarios are developed by embedding malicious intent gradually (jiang2024red, ), breaking down harmful prompts into benign sub-queries (yu2024cosafe, ; zhou2024speak, ; liu2024imposter, ), designing attack patterns (ren2024derail, ), and dynamically adjusting attack queries based on contextual feedback (li2024llm, ; yang2024chain, ; russinovich2024great, ). One contemporaneous multi-turn defense method (lu2025x, ) learns the safety boundary through fine-tuning without explicitly considering multi-turn context, and also cannot be used as a guard model across different LLMs. Current guard models typically introduce a separate model designed to moderate LLMs to filter out unsafe content (markov2023holistic, ; zeng2024shieldgemma, ; inan2023llama, ), but dynamically drifting context poses a challenge to existing reactive safety defense mechanisms. To the best of our knowledge, we are the first to safeguard LLM dialogues dynamically from jailbreaks through online filtering of harmful prompts before the prompts are sent to LLMs.
@@
In the literature of control and robotics, there is extensive research on learning-based controllers for dynamical systems that provide safety guarantees or certificates (boffi2021learning, ; herbert2021scalable, ; xiao2023barriernet, ; lindemann2021learning, ; chang2019neural, ; mazouz2022safety, ). Neural networks have been employed to parameterize control barrier functions (CBFs) to achieve forward invariance (robey2020learning, ; so2023train, ; zinage2023neural, ; dawson2022safe, ; dai2022learning, ): Once the system states enter the user-defined safe set, they remain within it indefinitely, thereby guaranteeing safety with neural certificates. Although neural CBFs can be successfully learned and verified for control-affined dynamical systems (chen2024learning, ; hu2024verification, ; mathiesen2022safety, ; wang2023simultaneous, ; rickard2025data, ), it is still challenging to guarantee the safety for non-analytical dynamical systems in latent space (hu2024real, ; wei2022safe, ; liu2023model, ; li2023system, ; shen2024bab, ; cheng2024robust, ), as the problem we tackle here. Recently, bajcsy2024human formulate a human-AI safety filter framework and miyaoka2024cbf introduce CBFs for LLM safety at the token level against single-turn jailbreaks. However, no existing work has explored dialogue-level safety for LLMs from the perspective of neural CBFs.
@@ 3 Problem Formulation
In this section, we formalize our approach to defending against multi-turn jailbreaks by enforcing invariant safety in the conversation with LLMs. Our framework models the conversation as an evolving dynamical system, where the hidden state represents the drifting context (e.g. the production process in Figure˜1), and each dialogue turn represents a state transition influenced by user queries and LLM responses in the language space. Given the language space $\mathcal{S}$}, at each turn $k$, LLMs receive a query $U_{k}\in{\mathcal{S}}$ from users and make a response $Z_{k}\in{\mathcal{S}}$ to that query, followed by the next round of user query $U_{k+1}\in{\mathcal{S}}$ and LLMs response $Z_{k+1}\in{\mathcal{S}}$. In accordance with the AI usage policies (openai2022usage, ), LLM responses $Z_{k},k-1,2,\dots,K$ should follow the AI safety rules and fall into the safe region specified by the user ${\mathcal{S}}_{0}$. However, in the multi-turn setting, the context drift of the query along the dynamic conversation may increase the vulnerability of the LLMs for jailbreaks, even though the response is safe at each turn. To this end, we introduce the concept invariant safety and associated measures to guarantee that once the LLM response falls into a safety region – which needs to be computed, the following LLM responses will stay within it no matter what future queries are given along the query context flow.
@@
For any safe but non-invariant responses $Z^{\prime}_{k}\in{\mathcal{S}}_{0}\setminus{\mathcal{S}}_{I}$ (e.g. the LLM response at turn 3 in Figure˜1), there exists a potentially harmful query $U^{\prime}_{k+1}\in{\mathcal{S}}_{context}^{(k)}$ such that the next LLM response $Z^{\prime}_{k+1}$ will inevitably go out of ${\mathcal{S}}_{0}$, resulting in LLM jailbreaks. Therefore, the safety invariance subset ${\mathcal{S}}_{I}$ is introduced to avoid non-invariant responses to achieve invariant safety against multi-turn jailbreaking attacks. In the following section, we first model the neural dialogue dynamics of multi-turn conversation with LLMs using the state-space representations. Then, we introduce the safety predictor based neural barrier function, followed by an invariant safety certificate learning framework.
@@ 4.1 Neural Dialogue Dynamics of Multi-turn Conversation with LLMs
Different from studying large language models (LLMs) dynamics of token generation in the literature (soatto2023taming, ; liu2024meanings, ; kong2024aligning, ; miyaoka2024cbf, ), we focus on the multi-turn human-AI interactive sentence-wise dynamics in the sentence embedding space and formulate it based on the state space representations in control theory. Given a $K$-turn conversation dialog with user query sentences $U_{k}\in{\mathcal{S}},k=1,\dots K,$ and response sentences of LLMs $Z_{k}\in{\mathcal{S}},k=1,\dots K,$ we first map them from the sentence language space $\mathcal{S}$} to the semantic and meaningful embedding space $\mathbb{R}^{n}$ using the pretrained sentence embedding model $f_{embedding}:{\mathcal{S}}\rightarrow\mathbb{R}^{n}$ (reimers2019sentence, ; fonseca2025safeguarding, ) as follows,
@@
We assume the dialogue with LLM is governed by the discrete-time state-transition equation parameterized by neural networks (NNs) $f_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ under initial state $\textbf{0}_{m}\in\mathbb{R}^{m}$, where the user query embedding $u_{k}$ serves as the control input at each turn $k$. However, the state representation $x_{k}$ is partially observable from the response embedding $z_{k}$ because the multi-turn dialogue dynamics is non-Markov due to the memory mechanism of LLMs. Therefore, we formulate the LLM response generation process through another NN-parameterized observation function $g_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where the LLM response embedding $z_{k}$ is the observed output from the hidden state $x_{k}\in\mathbb{R}^{m}$ and the user query embedding $u_{k}$. Assuming embeddings in Equation˜2 and states are consistent across different LLM seeds given fixed user queries and temperature, we have,
@@
To learn the state-space representation from the multi-turn dialogue, we construct the following mean square error (MSE) loss given $N$ trajectories of $K$-turn query and response embeddings $u^{(i)}_{k},z^{(i)}_{k}\in\mathbb{R}^{n},i=1,\dots,N,k=1,\dots,K$ from the pretrained embedding model in Equation˜2,
@@ 4.2 Neural Barrier Function based on Safety Predictor
On top of the language state-space dynamics above, we introduce the safety property in this section. Following the literature using LLM-based judge scores to evaluate the performance (ren2024derail, ; qi2023fine, ; zheng2023judging, ), we assume that for each conversation trajectory, the query and response embeddings $u_{k},z_{k}$ at $k$-th turn are associated with a discrete safety score $y_{k}\in{\mathcal{Y}}$ as a label from an LLM judge (qi2023fine, ). The scores of non-jailbreaking turns fall into the safe subset ${\mathcal{Y}}_{safe}\subset{\mathcal{Y}}$, which is equivalent to the user-specified safety region, i.e. $y_{k}\in{\mathcal{Y}}_{safe}\Leftrightarrow Z_{k}\in{\mathcal{S}}_{0}$. Using a more granular set of safety labels instead of simple binary ones allows for a more nuanced assessment of safety levels, providing richer information for training and evaluation. We adopt a NN parameterized safety predictor $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ to output the difference of predicted probability $p(\hat{y}_{k}\mid x_{k-1},u_{k})$ between the unsafe label and the most likely safe label. Inspired by cohen2019certified , the predictor $h$ is formally defined as,
@@
where the predicted label can be found through classification model $\hat{y}_{k}=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}p(y\mid x_{k-1},u_{k})$. It can be trained through the cross-entropy loss with $N$ trajectories of $K$-turn queries and responses and state-space dynamics in Equation˜3 ,
@@
Now we consider the state evolution in Equation˜3 during the multi-turn conversation by bridging the safe control theory (liu2014control, ; ames2014control, ) and the safety predictor. Since the user query sequences determine the state trajectory during the interactive conversation with LLMs, we first denote the query context embedding set at turn $k$ as ${\mathcal{U}}_{k}$, which evolves along the query context flow ${\mathcal{S}}^{(k)}_{context}$ due to multi-turn jailbreaking attacks. To prevent multi-turn jailbreaking, it suffices to ensure that the safety predictor always has safe predictions (negative outputs) given the previous state $x_{k}$ and all potential query embedding $u$ among query context embedding set ${\mathcal{U}}_{k}$ at each turn $k$. Formally, we define the neural barrier function (NBF) below as a safety index to show if the current state can be jailbroken or not along the conversation.
@@
Given the safety predictor $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ defined in Equation˜5, denote the query context embedding set at turn $k$ as ${\mathcal{U}}_{k-1}:=\{u\subset\mathbb{R}^{n}\mid u=f_{embedding}(U),\forall U\in{\mathcal{S}}_{context}^{(k-1)}\}$, and then the neural barrier function $\phi_{k}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ and the induced safe set ${\mathcal{X}}_{k}\subset\mathbb{R}^{m}$ are defined as,
@@
where state $x$ follows Equation˜3 and $\eta\geq 0$ is the steering threshold w.r.t the safe set ${\mathcal{X}}_{k}$.
Based on the NBF $\phi_{k}$, the safe set ${\mathcal{X}}_{k}$ is defined as the zero sublevel set of $\phi_{k}$ under the context embedding set ${\mathcal{U}}_{k-1}$ at each turn $k$. The larger the steering threshold $\eta$ is, the more shrinkage the induced safe set ${\mathcal{X}}_{k}$ will have. Therefore, $x_{k-1}\in{\mathcal{X}}_{k}$ indicates that the LLM cannot be jailbroken by any potential query $u$ given the conversation query context set ${\mathcal{U}}_{k-1}$, meaning the original safe response sentence $Z_{k-1}$ is within the safety invariance set, i.e., $Z_{k-1}\in{\mathcal{S}}_{I}$. The conversation in the language and embedding space is illustrated in Figure˜3.